### v3

# Thread: A Logic-Based Data Organization Paradigm for How-To Question Answering with Retrieval Augmented Generation

Abstract
Recent advances in retrieval-augmented generation (RAG) have substantially improved question-answering systems, particularly for factoid ‘5Ws’ questions. However, significant challenges remain when addressing ‘1H’ questions, specifically how-to questions, which are integral for decision-making and require dynamic, step-by-step responses. The key limitation lies in the prevalent data organization paradigm, chunk, which commonly divides documents into fixed-size segments, and disrupts the logical coherence and connections within the context. To address this, we propose Thread, a novel data organization paradigm enabling systems to handle how-to questions more effectively. Specifically, we introduce a new knowledge granularity, ‘logic unit’ (LU), where large language models transform documents into more structured and loosely interconnected LUs. Extensive experiments across both open-domain and industrial settings show that Thread outperforms existing paradigms significantly, improving the success rate of handling how-to questions by 21% to 33%. Additionally, Thread demonstrates high adaptability across diverse document formats, reducing retrieval information by up to 75% compared to chunk, and also shows better generalizability to ‘5Ws’ questions, such as multi-hop questions, outperforming other paradigms.
Thread: A Logic-Based Data Organization Paradigm for How-To Question Answering with Retrieval Augmented Generation
Kaikai An1,2, Fangkai Yang2, Liqun Li2, Junting Lu1,2, Sitao Cheng2,3, Shuzheng Si1, Lu Wang2, Pu Zhao2, Lele Cao2, Qingwei Lin2, Saravan Rajmohan2, Dongmei Zhang2, Baobao Chang1 1 Peking University 2 Microsoft 3 Nanjing University ankaikai@stu.pku.edu.cn,fangkaiyang@microsoft.com,chbb@pku.edu.cn thread.github.io
1 Introduction
Question answering (QA) is a foundational research topic in human-machine interaction (Allam and Haggag, 2012). Among the most advanced techniques, retrieval-augmented generation (RAG) enhances QA systems by organizing external documents into fixed-size chunks and retrieving relevant knowledge (Shao et al., 2023; Trivedi et al., 2023; Jiang et al., 2023; Asai et al., 2023). This routine is particularly effective in handling the ‘5Ws’ questions, such as ‘When is Shakespeare’s birthday?’, which typically require the data organization paradigm providing chunks containing the relevant knowledge, e.g. triples or documents about the topic entity (Yang et al., 2018; Jiang et al., 2019; Kwiatkowski et al., 2019; Stelmakh et al., 2022). However, the ‘1H’ questions, derived from Aristotle’s Nicomachean Ethics (‘how-to’ questions, Crisp 2014), remain largely underexplored [^1] The ‘5Ws’ represent What, Why, When, Where, and Who, and the ‘1H’ stands for How. These questions are in high demand in practical applications such as teaching us how to write code to achieve specific goals.
Central to problem-solving (Polya and Pólya, 2014) and human learning (Learn, 2000), how-to questions inherently involve complex processes that require interpretation and analysis (Deng et al., 2023b). For example, answering the question in Figure 1How to diagnose and fix a performance issue in a web application?’ involves a step-by-step decision-making process, i.e., first checking server load and response time, followed by optimizing server configuration based on user feedback. This dynamic, stepwise nature necessitates RAG systems to guide users through each step, adapting to specific contexts and providing precise and logical information. However, prevalent chunk-based data organization paradigm [^2] The term ‘chunk’ here refers to a general document splitting paradigm including chunks, sentences, phrases, etc. (Splitter, 2023; Chen et al., 2023; Gao et al., 2023), which divides documents into fixed segments, disrupts the logical coherence of content. As a result, RAG systems (Asai et al., 2023; Shao et al., 2023) struggle with how-to questions, often generating excessive, fragmented information that fails to maintain continuity between steps. To address this, a paradigm shift is needed, one that preserves the logical structure and stepwise nature of how-to questions.
Figure 1: An example of how-to questions with its decision-making process. We omit details such as the actions to check the server load and response time due to limited space.
In this paper, we propose Thread, a new logic-based data organization paradigm designed to handle how-to questions. The name Thread evokes the idea of ‘Pulling on the thread, the whole mystery started to unravel like a sweater.’ (Garcia and Stohl, 2011). Specifically, we introduce a new knowledge granularity named ‘logic unit’, comprising five key components and four different types (see §2.1 and §2.2). We employ a two-stage process (depicted in §2.3) to extract logic units (LUs) from documents. The first, optional stage is reformulating the original documents depending on their format and style, and the second focuses on extracting and merging LUs. In this way, Thread captures connections within the documents, breaking them into more structured and loosely interconnected logic units. When answering how-to questions, the system integrated with Thread enables a dynamic interaction manner. First, it retrieves relevant LUs based on their indexed headers. Then, the body of the selected LU provides the necessary content to generate responses for the current step. With user feedback, the linker in LU dynamically connects to other LUs, allowing the system to adapt its responses until the how-to question is comprehensively addressed.
We evaluate the effectiveness of Thread through experiments in two open-domain, Web Navigation (Deng et al., 2023a), Wikipedia Instructions (Koupaee and Wang, 2018), and one industrial setting, Incident Mitigation (Shetty et al., 2022). Experimental results demonstrate that while existing paradigms struggle with how-to questions, Thread excels at handling consecutive steps and consistently outperforms them, particularly in real-world incident mitigation scenario, with success rate improvements ranging from 21.05% to 33.33%. Additionally, Thread shows great superiority in processing diverse document formats, reducing both the number of retrieval units and the token length required for generation. Finally, we further validate the generalizability of Thread on open-domain tasks, where it outperforms the chunk-based paradigm in handling multi-hop questions. The main contributions of this paper include:
  • We highlight the challenges faced by current RAG systems in addressing how-to questions. To address the limitation of chunk-based paradigm, we propose Thread, a novel data organization paradigm that transforms original documents into structured, interconnected logic units.
  • Integrated with Thread, our system follows a dynamic interaction manner, guiding users incrementally through each step and adapting to their specific circumstances. Our system also brings more possibilities for an automation pipeline, solving how-to questions more efficiently.
  • Experimental results demonstrate that Thread significantly outperforms existing data organization paradigms across three scenarios. Furthermore, Thread efficiently handles various document formats, reducing the retrieval burden, and generalizes effectively to open-domain tasks.
2 Methodology
In this section, Thread leverages the internal logic and coherence of documents to construct the knowledge base, making it especially effective for addressing how-to questions. When converting documents into logic units, we classify how-to questions into two types: linear and dynamic. Linear how-to questions involve a fixed sequence of steps that do not require feedback or decision-making based on intermediate results. In contrast, dynamic how-to questions require decision-making at each step, with the process adapting dynamically based on previous actions, such as the example in Figure 1.
Below we first introduce ‘logic unit’ (LU) with its components and types, then explain how to extract LUs and construct the Thread knowledge base, followed by an illustration of how Thread integrates with existing systems.
2.1 Logic Unit: Retrieval Unit of Thread
Unlike the chunk-based paradigm, the logic unit consists of specially designed components that maintain the internal logic and coherence of a document, particularly when bridging consecutive steps in how-to questions. LUs serve as retrieval units that replace traditional chunks in RAG systems, ensuring a more structured and contextually relevant approach to information retrieval. Examples of LU in Table LABEL:tab:example_lu further illustrate such structure.
Prerequisite. The prerequisite component acts as an information supplement, providing the necessary context to understand the LU. For example, an LU may include domain-specific terminology such as entities or abbreviations. The prerequisite explains these terms and can generate new queries to retrieve LUs with more detailed information. Without this context, passing these LUs to an LLM-based generator could lead to hallucinations. Additionally, the prerequisite can function as an LU filter, containing constraints that must be met before the LU is considered in answer generation. This filtering ensures only
Header. The header summarizes the LU or describes the intention it aims to address, depending on the type of LU (refer to §2.2). For example, the header could be the name of a terminology if LU describes a terminology; if the LU describes actions to resolve a problem, the header describes the intent or the problem LU aims to resolve. Different from chunk that indexes the entire content, we use the header for indexing which serves as the key for retrieving the LU based on a query.
Body. The body contains detailed information on the LU, which is the core content fed into the LLM-based generator to generate answers. It includes specific actions or necessary information such as code blocks, detailed instructions, etc. This detailed content helps resolve the query mentioned in the header or provides a detailed explanation.
Linker. The linker acts as a bridge between logic units, enabling the dynamic process of how-to questions. Unlike the chunk-based paradigm, which relies on previous retrieval units that often lack direct clues, the linker in Thread provides necessary information to generate new queries for subsequent retrieval. Its format varies by LU type, serving as either a query for retrieving other LUs (“If-Then" condition) or an entity relationship. The edge of knowledge graph in traditional factoid questions is a special linker enabling navigation between related entities. When no further LUs are connected, the linker remains empty, isolating the current LU.
Meta Data. The meta data includes information about the source document from which the LU is extracted, such as the document title, ID, date, and other relevant details. This meta data is crucial for updating LUs when the source documentation is revised and reprocessed.
2.2 Logic Unit Type
When converting documents into logic units, Thread expands its scope beyond solving how-to questions, as documents often contain more than just solutions to these types of questions. Below are the common LU types identified in our experiments [^3] These LU types are summarized from our experiments with both industrial and public datasets. Additional LU types are provided in Appendix LABEL:appendix_lu_type, and more may emerge depending on the specific scenario., demonstrating Thread’s versatility in handling a wide range of document types:
Step. This is the most common type of LU for resolving how-to questions. Each LU body encapsulates detailed actions, including code snippets and resolution instructions. The LU prerequisite specifies the conditions or actions that must be completed prior to executing the current LU, serving a critical role in determining the appropriate entry point for a solution. For instance, when encountering a problem, multiple resolution paths may exist depending on the specific context. The prerequisite functions as a filtering mechanism during the LU selection process, eliminating candidates that do not satisfy the required conditions. Samely, the linker acts as a bridge between steps, directing the flow to subsequent LUs based on the outcome of the current step.
Terminology. This type provides detailed explanations of domain-specific terminology. For example, terms may share the same name or abbreviation in LU header but convey different meanings in LU body. The prerequisites for terminology LUs describe scenarios where the terminology typically appears and linkers are usually empty unless referencing extended terminology that depends on it.
Figure 2: The upper part shows the construction process of Thread, a two-stage process including reformulating documents into structured ones (a) and extracting and merging logic units (b). The bottom (c) illustrates the RAG system integrated with Thread. It retrieves relevant LUs based on query-Header similarity and filters out LUs that do not meet the current Prerequisites. The selected LUs are passed to LLMs to generate actions based on Body for execution. After execution, the Linker matches the results and generates a new query for the next retrieval iteration.
2.3 Thread: LU-based Knowledge Base
In practice, documentation is often unstructured and varies in format and style. Our approach to converting documentation into Thread involves a two-stage process to obtain LUs, as shown in the upper part of Figure 2.
Documentation Reformulation. This stage is optional, depending on the quality of the documentation. For example, in software engineering, Troubleshooting Guides (TSGs) often vary in format, include diverse types of information, and lack readability and detail, negatively impacting productivity and service health (Shetty et al., 2022). Due to such format, where some are clearly outlined and others are disordered, we avoid directly extracting LUs from original documents.
Instead, we first reformulate these documents into structured formats. By leveraging LLMs, we enhance the LLMs’ in-domain understanding by providing search capabilities and domain-specific context. This is followed by a refinement step to prevent overlooking details or hallucinating information. Figure 2 (a) shows the reformulation stage. But it is unnecessary for well-written documents like product help docs, which typically follow a linear how-to format. We instantiate this process by example in Appendix LABEL:appendix_instantiation.
LU Extraction and Merge. After reformulation, multiple LUs of varying types can be extracted from a single structured document (shown in Figure 2 (b)). Unlike chunk-based data organization commonly with fixed chunk sizes, LU granularity depends on content. For example, solutions to linear how-to questions typically form a single path from start to completion, with interconnected steps and no multiple execution outcomes encapsulated in one LU, such as an FAQ LU. However, for dynamic how-to questions with multiple possible outcomes, it is better to have one step per LU (Step LU), with Linkers navigating to the next LUs. Note that in dynamic how-to questions, not every step has multiple execution outcomes. If only one next step exists, the LUs can be merged to include both current and subsequent steps. Additionally, LUs with similar Headers and Bodies should be merged, extending the Prerequisite and Linker. More details about merging are provided in Appendix LABEL:tab:lu_merge.
LU update. In industry, documentation is often updated with each product version release. When this happens, we redo the above steps only for the updated documentation, identifying LUs in Thread with their Meta Data and replacing outdated LUs. As we extract and merge LUs, the collection of LUs from all documents forms the knowledge base, which serves as an essential component compatible with the current RAG system.
Dataset #Docs #Tasks #Steps #Chunks #LUs Dynamic Executable
Mind2Web 490 252 2094 6210 1089
WikiHow 97 97 2140 4225 774
IcM 56 95 323 413 378
Table 1: The statistics and characteristics of datasets, including the number of documents, LUs, etc.
2.4 Integrate Thread with QA System
To demonstrate how Thread works, we use dynamic how-to questions as an example. Figure 2 (c) shows the RAG system incorporating our Thread data organization paradigm.
LUs are indexed by their Headers. When an initial how-to question is submitted, the Retriever identifies the top-K most relevant LUs based on query-header similarity. The LLM-based Selector then checks the prerequisites of these LUs and filters out those that do not meet the current prerequisites, derived from the initial question or any available chat history. After selection, the body of the LUs is fed into the LLM-based Generator to produce an answer. If an execution engine is available, actions can be executed automatically; otherwise, the answer/action is presented to the user for manual execution. Once the action is executed, the Linker matches one of the possible outcomes and generates a new query for the next retrieval round.
Unlike traditional RAG systems, Thread-enabled systems can be potentially fully or semi-automated when integrated with execution engines. This integration offers greater automation and flexibility, as updating LUs automatically updates the system, compared to manually designed pipelines.
3 Experiments
3.1 Scenarios and Datasets
We evaluate Thread on two open-domain scenarios: Web Navigation, Wikipedia Instructions, and one industrial setting: Incident Mitigation.
Web Navigation. Mind2Web (Deng et al., 2023a) is a dataset designed for web agents to perform complex tasks on real-world websites based on language instructions. Each task is treated as a multi-choice question. At each step, the input consists of HTML code, an instruction, and a set of choices, while the output is the selected choice, operation, and an optional value. We treat each one as a ‘dynamic how-to question’, with multiple possible outcomes depending on executed actions.
Wikipedia Instructions. WikiHow (Koupaee and Wang, 2018) [^4] https://www.wikihow.com provides step-by-step procedural guides, each titled ‘How to’, with a fixed sequence of steps. This dataset is used to evaluate Thread on linear how-to questions.
Incident Mitigation. IcM (Shetty et al., 2022; An et al., 2024) is critical for managing large-scale cloud services, where engineers rely on Troubleshooting Guides (TSGs) to resolve incidents. Each step can lead to different outcomes based on system states, making this dataset suitable for testing Thread on dynamic how-to questions. Unlike the open-domain datasets, we conduct a human evaluation with twenty on-call engineers (OCEs) responsible for incident mitigation. Each OCE mitigates five incidents, with baseline methods randomly assigned per incident to prevent familiarity bias. Our RAG system attempts automated mitigation, and if a failure occurs, an OCE intervenes before the system resumes [^5] Varying across new-hire and experienced OCEs. Note that one OCE’s data was contaminated during the experiment, so we removed that OCE’s data..
We collect documents from open sources, enterprises, and LLMs across scenarios. Appendix A.1 outlines the collection process, Appendix A.2 provides dataset examples, and Table 1 summarizes dataset statistics.
3.2 Baselines and Metrics
Previous work on the Mind2Web dataset, such as MINDACT (Deng et al., 2023a), has not approached it as how-to questions, instead using In-Context Learning (ICL) or Supervised Learning (SL). For fair comparison with our LLM endpoints, [^6] We use GPT-3.5 and GPT-4 (version 1106-preview). we re-implement MINDACT using the same demonstrations and include chat history as context. In this work, we treat Mind2Web as dynamic how-to questions and address them with our RAG system (Appendix LABEL:appendix_rag shows how we adapt Mind2Web into how-to questions.). We compare Thread against document-based and chunk-based data organization paradigms, also used as baselines for the WikiHow and IcM datasets. Details on evaluation metrics, experimental setup (prompts, etc.), and data organization paradigms are provided in Appendix LABEL:appendix_metrics, LABEL:appendix_exp, and LABEL:appendix_paradigms, respectively.
Mind2Web Cross-Task
Method Model Paradigm Ele. Acc Op. F1 Step SR SR
w/ GPT-3.5 ICL 40.69 49.66 33.91 1.59
w/GPT-4 ICL 62.80 60.37 51.81 10.32
MINDACT (2023a)  w/ Flan-T5XL* SL 55.10 75.70 52.00 5.20
w/ GPT-4 Chunk 64.23 65.96 58.45 8.73
w/ GPT-4 Doc 63.80 65.89 58.36 11.51
RAG w/ GPT-4 Thread 68.29 69.53 61.94 12.30
Table 2: Experimental results on Mind2Web. ‘*’ represents taking results from the original paper. Deng et al. (2023a) formulate web navigation as a series of multiple-choice steps, where each step requires selecting the correct HTML code and actions. SR refers to the overall Success Rate.
3.3 Main Results
Web Navigation
Table 2 presents the overall performance on Mind2Web, comparing our method with baselines and both doc-based and chunk-based RAG methods. We find that incorporating informative documents, regardless of the data organization paradigm, significantly improves performance. RAG methods outperform both ICL and SL approaches, with the doc-based RAG method achieving results comparable to MINDACT’s best performance. Notably, Thread surpasses all RAG baselines, improving Ele. Acc by 4.06%, Step SR by 3.49%, and SR by 3.57%. MINDACT-SL achieves the highest Op. F1 due to label imbalance [^7] Predicting all operations as “CLICK” results in an Op. F1 of 79.90%., biasing the model toward frequent operations.
Incident Mitigation
Paradigm SR Step SR P.F. Step SR HI Turns
40.51 60.90 60.90 30.10 3.14
Chunk 28.95 53.16 43.05 46.84 6.84
43.86 63.90 63.90 36.09 2.98
Doc 31.58 57.89 42.11 42.11 6.53
77.19 88.72 84.21 11.28 2.56
Thread 52.63 84.21 68.95 15.79 5.74
Table 3: Experimental results on Incident Mitigation. Incidents are divided into simple and hard groups. P.F. Step SR: the perctenage of successful steps before the first failure. HI: the steps requiring human intervention.
Incident Mitigation
Table 3 highlights the advantages of Thread in handling complex dynamic how-to questions. Both chunk-based and doc-based RAG methods struggle with incident mitigation, as their low SR and P.F. Step SR scores indicate difficulty in connecting subsequent steps based on the current state. In contrast, Thread achieves the highest performance across all metrics, notably improving SR from 21.02% to 33.33%. More importantly, its superior P.F. Step SR demonstrates its ability to dynamically link steps based on user feedback, reducing human intervention. As a result, Thread not only achieves the highest Step SR but also minimizes interaction turns, effectively mitigating both simple and complex incidents.
Wikipedia Instructions
Table 4 compares different data organization paradigms on WikiHow under two settings. For single-turn, where the RAG system generates the entire plan in one step, Thread outperforms the doc-based by 5.15%, providing more concise and effective information. In contrast, the chunk-based achieves a significantly lower SR of 19.59%, highlighting the challenge of retrieving relevant chunks without maintaining the document’s logical flow. For multi-turn, iterative retrieval improves performance over single-turn, demonstrating the benefits of a step-by-step approach for how-to questions. The SR increases from 58.76% to 68.04% for the doc-based and from 63.91% to 72.16% for Thread. As in single-turn case, the chunk-based method disrupts internal logic, resulting in a low SR of 20.62%. Overall, Thread achieves the highest SR of 72.16%, emphasizing its effectiveness in preserving and modeling step dependencies. We also extend our experiments to LLaMA3-70B to show the generalibity of Thread in Appendix LABEL:sec:backbone.
WikiHow
Paradigm SR Precision Recall F1
Single-Turn
Chunk 19.59 60.20 25.16 35.49
Doc 58.76 77.71 57.93 66.37
Thread 63.91 83.43 71.48 76.99
Multi-Turn
Chunk 20.62 52.95 25.54 34.45
Doc 68.04 87.10 70.65 78.02
Thread 72.16 89.77 73.36 80.74
Table 4: Experimental results on WikiHow with different interaction manners and paradigms.
4 Analysis
4.1 Ablation on RAG System Settings
This section presents an ablation study on key components of our RAG system, using the Mind2Web dataset. While retriever and generator variants have been explored in prior work (Gao et al., 2023), we utilize text-embedding-ada-002 (OpenAI, 2022) for retrieval and GPT-4 as the generator.
Paradigm Ele. Acc Op. F1 Step SR
ICL 62.80 60.37 51.81
$w/o.$ historical steps 56.97 59.09 50.38
Chunk 59.94 62.58 54.39
$w/o.$ chunk selection 64.23 65.96 58.45
Thread 68.29 69.53 61.94
$w/o.$ LU selection 67.05 68.43 60.79
Table 5: Ablation study of integrating chat history and retrieval unit selection on Mind2Web.
Multi-turn Interaction. As shown in Table 4, the multi-turn setting outperforms the single-turn setting in answering how-to questions. We adopt the multi-turn setting for all scenarios.
Chat History. Chat history enables the system to reference previous actions and results, improving decision-making. Table 5 shows that excluding chat history results in performance degradation. Therefore, we include chat history for all RAG-based methods in our experiments.
Retrieval Units Selector. As described in §2.4, the Selector identifies the most relevant retrieval units from the top-K retrieved units. Table 5 illustrates the effect of the retrieval unit selector. Without the selector, performance drops by 4.29% in Ele. Acc and 3.38% in Op. F1. The selector improves all metrics when applied to Thread. Unlike LU selection, which filters irrelevant LUs based on prerequisites, chunk selection ignores inter-chunk connections and may exclude relevant chunks. Thus, we activate the selector only for Thread.
4.2 Comparison with Different Paradigms
We compare several RAG data organization paradigms (details in Appendix LABEL:appendix_paradigms), including Semantic Chunking (Kamradt, 2024), Proposition (Chen et al., 2023), recursive chunking (Splitter, 2023) (chunk-based), entire document (doc-based), and GraphRAG (Edge et al., 2024). As shown in Table 6, Thread outperforms all baselines across metrics. While Semantic and Proposition merge semantically similar sentences using LLMs, they fail to capture logical relationships. And GraphRAG underperforms on how-to questions due to its focus on entity-level relations over document logic. In contrast, Thread retrieves smaller, logic-driven units, yielding higher efficiency and accuracy. Appendix LABEL:case_study illustrates the difference when handling the question in Table LABEL:tab:example_lu between GraphRAG and Thread, with GraphRAG producing flawed outputs. Appendix LABEL:appendix_result analyzes Thread ’s cost and scalability, confirming its high performance at acceptable costs and potential for large-scale datasets.
Paradigm Ele. Acc Op. F1 Step SR #Tokens in RU
Doc 63.80 65.89 58.36 663.84
Recursive 64.23 65.96 58.45 695.77
Semantic 65.14 67.30 59.93 1337.16
Proposition 62.37 64.78 56.78 790.14
GraphRAG 63.20 65.09 56.90 630.75
Thread w/o. 67.05 68.43 60.79 772.67
Thread 68.29 69.53 61.94 157.10
Table 6: Analysis of data organization paradigms.
4.3 Superiority over Different Doc Formats
As real-world documents vary in format, we test our LU extraction method on diverse structures, including structured markdown, hierarchical guidelines, tabular checklists, and narrative documents (details in Appendix LABEL:appendix_mind2web). Table 7 demonstrates that Thread consistently outperforms the chunk-based paradigm across all metrics, improving Ele. Acc by up to 9.61%, Op. F1 by up to 8.43%, and Step SR by up to 8.51%. These results underscore Thread’s ability to effectively handle a variety of real-world document resources. Notably, Thread achieves the highest performance with structured documents, which facilitate the creation of a higher-quality knowledge base.
Format Paradigm Ele. Acc Op. F1 Step SR
Chunk 64.23 65.96 58.45
Structured Thread 68.29 69.53 61.94
Chunk 60.60 63.46 55.06
Hierarchical Thread 66.57 67.89 60.08
Chunk 56.30 59.26 51.43
Tabular Thread 65.71 67.69 59.55
Chunk 56.63 60.39 51.66
Narrative Thread 66.24 68.22 60.17
Table 7: Analysis of different document formats.
Figure 3: Experimental results on the generalization of Thread to 2WikiMultiHopQA.
4.4 Generalization to Open Domain Questions
To explore whether Thread can be extended to open-domain questions, we evaluate Thread on multi-hop questions, 2WikiMultiHopQA (Ho et al., 2020). Specifically, we sample 200 questions of two types, ‘inference’ and ‘compositional’ type, like “Who is the maternal grandfather of Abraham Lincoln?” and “Who is the founder of the company that distributed La La Land film?” In our experiment, we use the relations between entities as ‘Linker’. The findings from Figure 3 demonstrate that THRED largely surpasses the chunk-based paradigm on multi-hop questions, validating the generalization of Thread on diverse questions.
5 Related Work
Data Organization Paradigm in RAG
The data organization process is a critical pre-stage of RAG methods where documents are segmented following certain data organization paradigms. The most common data organization paradigm is splitting documents into retrieval units (Gao et al., 2023). These retrieval units vary in granularity such as phrases, sentences, propositions (Chen et al., 2023), chunks (Kamradt, 2024), etc. Coarser-grained units contain more information but introduce redundant noise, while finer-grained units have lower semantic integrity and often require retrieving more units to gather comprehensive information. However, the chunk-based data organization paradigm ignores the logical and relational connections between chunks, potentially disrupting the inherent logic flow in documents. Another paradigm constructs documents into knowledge graphs (KG), where retrieval units include entities, triplets, etc. (Gaur et al., 2022; Sen et al., 2023; He et al., 2024; Wang et al., 2024). While these approaches emphasize semantic and lexical similarities between retrieval units, their effectiveness on how-to questions remains limited. This is largely due to the inherent difficulty in capturing the logical structure required by such questions. In particular, current methods such as GraphRAGEdge et al. (2024), primarily model relationships between different chunks, which fails to account for the more complex, logic-driven connections between chunks that go beyond entity-level similarity.
Information Retrieved by RAG
The effectiveness of RAG methods depends on the generator’s ability to utilize retrieved information and the quality and quantity of that information. Insufficient question-relevant information can cause hallucinations in LLM-based generators (Li et al., 2023; Zhang et al., 2023), making it crucial to improve the retrieval process. Traditional one-round retrieval methods (Guu et al., 2020; Lewis et al., 2020) often fail to gather all necessary information due to their reliance on the similarity between query and retrieval units (Gan et al., 2024). Advanced RAG methods use query rewriting and expansion (Shao et al., 2023; Trivedi et al., 2023; Kim et al., 2023) or iterative retrieval (Shao et al., 2023; Jiang et al., 2023; Asai et al., 2023) to collect more information. However, these approaches still struggle with how-to questions, which require making next-step decisions based on the current retrieved units, unless the current retrieved units contain clues that lead to the next step. The main issue is the lack of connections between retrieval units, which prevents effective retrieval and the gathering of sufficient information.
6 Conclusion
In this paper, we address the overlooked category of handling how-to questions in QA systems by proposing Thread, a novel data organization paradigm that captures logical connections within documents. By introducing a new knowledge granularity called ‘logic unit’, Thread restructures documents into interconnected logic units that are compatible with RAG methods. Extensive experiments show that Thread significantly outperforms existing paradigms, improving performance while reducing the knowledge base size and minimizing the information needed for generation.
Limitations
This work focuses on evaluating the effectiveness of Thread by designing how-to questions in three specific scenarios, covering both linear and dynamic how-to questions. However, several limitations point to future directions. First, as our logic-based knowledge base can coexist with the original chunk-based knowledge base, we only extend our method to multi-hop questions, more open-domain tasks should be considered. Next, while extracting logic units involves an initial cost in terms of LLM usage, this is a one-time process. Once the knowledge base is constructed, it provides significant advantages for industrial applications, particularly in terms of subsequent updates and maintenance.
Acknowledgement
We thank all reviewers for their great efforts. This work is supported by the National Science Foundation of China under Grant No.61876004, which covered personal research efforts, conference registration fees of K.K. An, S.Z. Si and B.B. Chang.
Appendix A Scenarios and Datasets
A.1 Documents under Each Scenario
As the Mind2Web dataset lacks relevant documents, we create retrieval documents tailored to the RAG system. Assuming that each website has help docs applicable across different tasks, we use the "Cross-Task" test set, selecting examples from the training set to craft informative documents for each website. And we follow Wang et al. (2023) to brainstorm different formats of documents. More details about collecting documents are shown in Appendix LABEL:appendix_mind2web.
For the WikiHow dataset, we utilize publicly available Windows Office Support Docs [^8] https://github.com/MicrosoftDocs/OfficeDocs-Support as retrieval documents. We select around 100 tasks from WikiHow tagged with Microsoft products like Word, PowerPoint, and Teams, each containing 10 to 40 steps related to Windows operations.
For the IcM dataset, we collect 56 TSGs from an enterprise-level engineering team responsible for a large-scale cloud platform. The selected incidents in the IcM dataset can be resolved using the knowledge provided in these TSGs.
A.2 Example of Each Dataset
We list one example of each dataset in Table LABEL:tab:example_dataset, and we distinguish how-to questions into ‘linear’ and ‘dynamic’ types, where linear how-to questions often involve a fixed sequence of steps that do not require feedback or decision points based on intermediate outcomes.
Dataset Example
Mind2web
<html> ... </html> Based on the HTML webpage above, try to complete the following task Task: Book the lowest-priced and quickest flight for 5 adults and 1 child on May 20 from Mumbai to any airport near Washington. Previous actions: None What should be the next action? Please select from the following choices (If the correct action is not in the page above, please select A. None of the above’): A. None of the above B. <div id=0> <input radio triptype roundtrip true /> <label> <span> C. <label id=1> <span> Search flights one way </span> <span> One D. <a id=2> <h3> Celebrate World Wish Day </h3> <p> Support E. <h2 id=3> Help </h2> F. <a id=4> <img follow us on twitter - /> </a> C. Action: CLICK WikiHow
"Problem": "How to Add Captions to Tables in Microsoft Word", "Solution Steps": [ "Select the table to which you want to add a caption.", "Using your mouse, click and drag over the entire table to select it.", "Right-click (or ctrl-click) the table and select Insert Caption.", "Enter your caption.", "Type the caption for this table into the \"Caption\" field.", "Select a caption label.", "Customize your caption numbers (optional).", "Choose where to place your caption.", "Click the \"Position\" drop-down menu, and choose whether to place the caption above or below the table.", "Click OK to add your caption to the table.", "Format your captions." ]
IcM